A fourth grade class decides to enclose a rectangular garden using the side of the school as one side of the rectangle. Suppose they have 32 feet of fence to enclose the garden. Express the are of the garden as a function of x. What should the dimensions of the garden be in order to maximize the area?
Answer by checkley77(12844)

This is one of the tricky ones.
One would think the garden would be a square BUT it isn't
First we divide the fencing by 4.
32/4=8 ft. this is the width.
2 times the width=2*8=16 for the length
So we have a perimeter of:
8+16+8=32 ft. for the perimeter.
Now for the area.
8*16
128 ft^2. for the maximum area.
Proof:
Try the sides of:
7.9+16.2+7.9=32
Area=7.9*16.2=127.98 ft^2
Try the sides of:
8.1+15.8+8.1=32
area=8.1*15.8=127.98 ft^2